Determine the derivative of the function "f(x) = x^2sinx" at "x = \u03c0\/2"


1
Expert's answer
2021-06-07T16:59:38-0400
"f'(x)=2x\sin x+x^2\cos x"

"f'(\dfrac{\pi}{2})=2(\dfrac{\pi}{2})\sin (\dfrac{\pi}{2})+(\dfrac{\pi}{2})^2\cos (\dfrac{\pi}{2})=\pi"

"f'(\dfrac{\pi}{2})=\pi"